Algebra 1: the length of a rectangular park is 3 times the width.if the area of the park is 300 square meters. Then what is the length and with of the park and the equation
Answer by checkley79(3341)

AREA=LW
L=3W
300=2*3W*W
300=3W^2
W^2=300'3
W=SQRT100
W=10 ANS. FOR THE WIDTH.
L=3*10=30 ANS. FOR THE LENGTH.
PROOF:
300=30*3
300=300